OpenStudy (anonymous):

1. Find f(4) for the following function. f(x) = 2.9x - 5 16.6 -6.6 -16.6 6.6

jimthompson5910 (jim_thompson5910):

f(x) = 2.9x - 5 f(4) = 2.9(4) - 5 ... replace every x with 4 f(4) = ???

OpenStudy (anonymous):

I really don't get it Lol

OpenStudy (anonymous):

u only replace the second x the first fx stand for the function of x

OpenStudy (anonymous):

f(x)=2.9x-5 f(4)=2.9*4-5=

OpenStudy (anonymous):

f(4)=11.6-5=6.6

jimthompson5910 (jim_thompson5910):

2.9(4) means 2.9 times 4

jimthompson5910 (jim_thompson5910):

so you just need to evaluate 2.9*4 - 5

OpenStudy (anonymous):

Ohh omg im so dumb Lmao thanks!! :)